# Learning Machine Morality through Experience and Interaction

### v2

@@ 3.2 Case Study: Morally-constrained Reinforcement Learning
Taking inspiration from the full top-down tradition, researchers have used logic to define constraints for training RL agents. For example, ? (?) use defeasible deontic logic to implement a normative supervisor onto a learning agent. This supervisor takes in a proposed input action set, applies a set of norms, and outputs a reduced set of actions for a Q-learning agent to choose from - this can either be a strictly ‘legal’ set of actions, or a relaxed set of ‘permissible’ actions if no fully satisfactory options are available in the strict set. The ‘ethical’ dimension in that work is implemented as a centralized mechanism or supervisor overseeing a learning agent, and - as discussed - the normative constraints are hard-defined, which makes this the most top-down of our case studies.
@@ 3.3 Case Study: Constitutional AI
As a second case study, we present the example of Constitutional AI, developed and implemented by ? (?). This approach is specific to the type of systems known as Large Language Models (LLMs), and ethically-informed RL is used here at the fine-tuning stage, i.e., after a model has been pre-trained on a large amount of linguistic data. As such, the end product of this process is not an agent but a predictive system; however, as we discuss below, the ethical fine-tuning step of the training can be perceived as agent-based, so still has relevance to the present discussion.
@@
Constitutional AI relies on one LLM providing feedback to some LLM being currently trained (either itself or an external one). The constitutional methodology has been proposed as a potentially safer, more controlled and interpretable alternative to the fully bottom-up RLHF approach used in other LLMs (?). The feedback model represents a ‘constitution’ of different principles, presented sequentially and defined via explicit prompts (e.g., ‘Please choose the assistant response that’s more ethical and moral. Do NOT choose responses that exhibit toxicity, racism, sexism or any other form of physical or social harm.’). The principles themselves are based on a combination of human defined preferences such as the UN Declaration of Human Rights, certain digital companies’ terms of service (to reflect the more recent digital dimensions of safety), and a set of other preferences defined by a team of researchers behind Constitutional AI (?) and others (?).
@@
Once a target LLM receives feedback based on these principles, its weights are fine-tuned with RL using the feedback as a reward signal. Thus, the model is fine-tuned to be more likely to produce outputs which would be considered safe by a constitution of potential ‘critic’ models with diverse preferences. An extension of this approach based on crowd-sourced constitutional principles is called Collective Constitutional AI (?) and may prove promising in the future in generating more generally aligned agents.
@@
In summary, according to this methodology the moral principles are explicitly defined via prompts, but the mapping between those principles and a change in behavior (i.e., the change in language model outputs) is based on a learned representation which is very difficult for humans to interpret, which poses a disadvantage in terms of safety. Furthermore, while the approach is certainly promising, it must be noted that ? (?) only managed to use the constitutional principle for improving a model’s harmlessness - it remains to be seen how successful the methodology would be at promoting other values such as helpfulness.
3.4 Case Study: Social Dilemmas
@@
Finally, we consider a case study that combines top-down principles and learning in a more equal weighting. This case study relies on the domain of two-player general-sum games known as social dilemmas (?, ?). Though these games constitute stark simplifications of reality, they are intended to model and reflect the general underlying structure of everyday social situations, including those involving multiple AI systems or agents.
@@ 3.4.1 Social Dilemma Matrix Games
Social dilemma games simulate situations in which each agent make decisions facing a trade-off between individual interest and societal benefit. A widely studied subset of these games is the two-player, two-action matrix game, in which a pair of players simultaneously choose an action (e.g., Cooperate or Defect), without the opportunity to communicate, each receive a certain payoff which depends on the pair of actions chosen.Higher payoff here represent the fact that that outcome is preferred by the player. What makes it a dilemma is the fact that players maximizing their individual payoffs are likely to learn policies that end up being sub-optimal in terms of long-term individual and social outcomes.
Three classic matrix dilemma games from Economics and Philosophy that are relevant to moral choice are the Prisoner’s Dilemma (?), Volunteer’s Dilemma (or ‘Chicken’) (?), and Stag Hunt (?; see payoffs for row vs. column player in Table 2). In the Prisoner’s Dilemma, mutual cooperation would achieve a Pareto-optimal outcome (in which one player cannot be made better off without disadvantaging the other) - but each individual player’s best response is to defect due to greed (facing a cooperator, they benefit from defecting and getting the highest payoff $T$) and fear (facing a defector, they suffer by cooperating and getting the lowest payoff $S$). In the one-shot game, this reasoning applied by both players leads to mutual defection and thus a sub-optimal outcome for both (e.g., the IPD game in Table 2 demonstrates that both players would have preferred mutual cooperation with payoff $3$ over mutual defection with payoff $2$). In the Volunteer’s Dilemma, a selfish or rational player may choose to defect due to greed (i.e., not volunteer in hope that someone else does), but if both do so, both obtain the lowest possible payoffs (i.e., no one volunteers, and the society suffers). Finally, in the Stag Hunt game, two players can cooperate in hunting a stag and thus obtain the greatest possible payoff each; however, given a lack of trust between the players (i.e., each player fears a non-cooperative partner), either may be tempted to defect and hunt a hare on their own instead, decreasing both players’ payoffs as a result.
@@
Thus, in all three games the best-response (i.e., ‘rational’) strategies are not the ones leading to the better mutual outcome. ? (?) argued that repeated versions of these dilemma games in particular are relevant to morality, since, with repeated interaction, complex sets of strategies can evolve - for example, involving actions to punish one’s opponents for past wrongdoing, or to influence their future actions. Repeated versions of the games can be implemented as learning environments for RL agents, as we discuss in Section 3.4.2. Due to instabilities in the multi-agent environment, calculating predicted equilibria in these situations is not always computationally feasible, so simulation methods are required in order to study potential emergent behaviors and outcomes (?, ?).
Other types of games in which it is possible to model human morality are sequential games (?, ?, ?), games with more than two players (?, ?), and games with incomplete information, first introduced by ? (?). We provide these in Appendix B for context, but focus on the simpler and more abstract two-player matrix games in the subsequent formalization due to their generalizability to a wide range of situations.
@@ 3.4.2 Reinforcement Learning in Social Dilemmas
A social dilemma environment can be implemented for RL agents as a two-player iterated game, played over $T$ iterations (?). At each iteration, a moral player $M$ and an opponent $O$ play the one-shot matrix game corresponding to a classic social dilemma (see Table 2 for examples). At every time step $t$, the learning agent $M$ observes a state, which is the pair of actions played by $O$ and $M$ at the previous time step $t-1$: $s_{M}^{t}=(a_{O}^{t-1},a_{M}^{t-1})$, and chooses an action (Cooperate or Defect) simultaneously with the opponent $O$: $a^{t}_{M},a^{t}_{O}\in\{C,D\}$. The player $M$ then receives a reward $R_{M}^{t+1}$ and observes a new state $s_{M}^{t+1}$. In traditional RL, this reward would be provided by the environment (i.e., extrinsic reward), and in the social dilemma games it would correspond directly to the payoffs from the game.
@@ 3.4.3 Moral Preferences as Intrinsic Rewards
? (?) and ? (?) draw interesting parallels between intrinsic motivation and rewards. The implication for RL is that effective learning can result from re-defining an agent to follow some internal reward signal rather than only rewards from the environment. Early successes with this approach were reported by ? (?) in curiosity-driven exploration, where curiosity is modeled as an intrinsic reward obtained when the agent visits unfamiliar states.
@@
The intrinsic rewards framework is directly applicable to the idea of learning morality from experience. A small set of existing works in Multi-Agent RL modified the reward signal for agents in social dilemma games to embed an internal moral value or preference into reward functions (?, ?, ?, ?). In these works, moral preferences have either been expressed as an operation on the extrinsic reward, a combination of the player’s own reward and that of their opponent(s), or a conditional function based on actions performed previously. A related line of work modifies reward functions to include an awareness of the opponents’ learning (?, ?) - but these lie beyond the scope of the studies reviewed here.
@@
A visualization of the intrinsic rewards approach is presented in Figure 3 - using the terminology defined in Section 3.1 and the two-player social dilemma games as a case study. In this formulation, for player $M$, we contrast traditional agents that learn according to an extrinsic game reward $R_{M_{extr}}$, which is simply the payoff associated with the joint actions $a_{M}^{t},a_{O}^{t}$ (as defined in Table 2), versus agents that learn according to an intrinsic reward $R_{M_{intr}}$, based on some moral principle. In Section 3.4.4, we expand on the types of principles that can be encoded as intrinsic rewards.
@@
Figure 3: A step in the RL process, from the point of view of a learning moral agent $M$ playing against a learning opponent $O$. Agent $M$ calculates their moral (intrinsic) reward according to their own ethical framework, which may or may not consider the game (extrinsic) reward coming from playing the dilemma.
@@
In practice, both extrinsic and intrinsic rewards can be combined in a multi-objective manner (?) to create an agent able to pursue multiple goals (for example, trying to obtain both moral reward and the reward from the task at hand - ?). This was the approach used by ? (?) to create their ‘curious’ yet task-specific agent. Furthermore, multi-objective RL approaches can also be used to combine multiple intrinsic moral rewards to create a multi-faceted moral agent, which allows for the creation of more complex and comprehensive moral agents.
@@
We implement a decentralized RL approach where each agent chooses an action in the game without a central planner. A parallel line of work exists in the modeling and ethics literature, where an ‘ethical’ agent is implemented as a centralized mechanism / supervisor (e.g., ?) - this has applications in entirely different domains such as governmental policy design (?), and lies beyond the scope of this work.
3.4.4 Formalizing Existing Ethical Frameworks as Rewards for AI Agents
@@
A variety of moral frameworks has been developed over the millennia in the fields of Moral Philosophy, Psychology, Biology and Economics. We review a representative set of these in Table 3, formulating them as potential reward functions for agents in two-player games. A growing body of RL work demonstrates that mathematical formulations of these frameworks are able to effectively support more prosocial and cooperative behavior in societies of agents (?, ?, ?).
@@
Utilitarian or Altruistic or Spiteful ?, ?, ?, ?, ? Sum of all players’ payoffs, with various weights on on own vs. others’ payoffs. either Consequentialist
@@
Deontological ?, ?, ?, ?, ? Positive (/negative) reward if the action taken adheres (/violates) to a pre-defined norm. Agent may enforce norm in others or not. External Norm-based
@@
Virtue Ethics (can be based on the above) ?, ?, ? Reward for adhering to a single specific virtue (e.g., cooperativeness), or a mix of virtues (e.g., sum cooperativeness + equality rewards). Internal either or both
@@
Table 3: Moral frameworks proposed in past literature in Philosophy, Psychology, Biology and Economics. The table also specifies the type(s) of rewards that can be used for their definition in a RL framework.
@@
We classify possible moral agent types in Table 3 into Consequentialist or Norm-based. Consequentialist morality focuses on the outcomes or consequences of an action. In dilemma games, this can be easily implemented as an operation on the two players’ payoffs. For example, this could include objective functions that are Utilitarian (?), defining actions as moral if they maximize total utility for all agents in a society, or Altruistic or Spiteful (?), where an agent is inclined to maximize or minimize their opponent’s winnings, perhaps without even considering their own. Consequentialist morality also includes fairness-based objective functions. These can be based on the difference between two players’ rewards, including the inequity aversion models of ? (?, ?), or other calculations of the ratio between payoffs with different preferences for fair or equal outcomes (?, ?, ?). Finally, consequentialism also accounts for the principles of distributive justice of ? (?), in which agents distribute resources without knowing where they will end up in that distribution.
@@
Norm-based morality, on the other hand, focuses on defining actions as moral so long as they adhere to a society’s external norms or to the specific agent’s duties, regardless of the consequences further down the line. This framework includes the Deontological ethics of ? (?), which identifies specific acts as morally required, permitted or forbidden. An example norm for dilemma game-playing agents may be conditional cooperation, in which an agent is expected to cooperate against a cooperative opponent (?). Further norm-based objective functions have been formalized by ? (?, ?, ?, ?), who propose various weights on personal versus social importance of the norm at hand.
@@
A third type of ethical framework that is often distinguished is Virtue Ethics of ? (?) - in this line of reasoning, moral agents act according to their certain internal virtues. In practice, these virtues themselves often have Consequentialist or Norm-based foundations - consider, for example, the virtues of fairness or care, respectively (?, ?). The key distinction of Virtue Ethics is that a single agent may rely on more than one type of virtue, and different agents may weigh the virtues differently against one another (?, ?), so a more expressive way of modeling virtue ethics might be through a weighted, multi-objective paradigm - such as that suggested by ? (?).
@@
In Table 3, we additionally distinguish agents by the external or internal nature of their moral motivation - while each of these frameworks can be formalized as an intrinsic reward function for the RL agent, sometimes the moral aspect of the intrinsic motivation comes from outside the agent themselves. An example of this is where certain moral norms exist externally in a society, so the agent prefers to adhere to them in their individual behavior.
@@
The final step for formalizing morality as intrinsic motivation for RL agents, then, is to formulate precise mathematical definitions of rewards based on these frameworks. Using the domain of social dilemmas introduced above, ? (?) propose a formulation in terms of actions and/or outcomes in a two-player, two-action social dilemma game. For illustration, we present their mathematical definitions of moral rewards at a certain time step $t$ in Table 4 (?).
@@
Table 4: Definitions of a sub-set of intrinsic moral reward types, from the point of view of the moral agent $M$ playing versus an opponent $O$.
@@
Norm-based morality, on the other hand, is operationalized in terms of positive reward for an agent whose action $a^{t}_{M}$ adheres to a given norm, or negative reward for its violation. For example, a Deontological agent following the norm of conditional cooperation (defined above) receives a negative reward of value $\xi$ for defecting against an opponent who previously cooperated. The other norm-based agent Virtue-kindness follows a simple cooperative norm, obtaining a positive reward $\xi$ for cooperating, regardless of what their opponent did previously.
Finally, these examples also outline one multi-objective reward function Virtue-mixed, which linearly combines the two separate virtues ‘equality’ and ‘kindness’, with a weighting parameter $\beta$, and a normalized parameter $\hat{\xi}$ used to fit the ‘kindness’ element of the reward to the same scale as equality.

### v3

@@ 3.2 Case Study: Morally-constrained Reinforcement Learning
Taking inspiration from the full top-down tradition, researchers have used logic to define constraints for training RL agents. For example, ? (?) use defeasible deontic logic to implement a normative supervisor onto a learning agent. This supervisor takes in a proposed input action set, checks these for compliance against a set of norms, and outputs a reduced set of actions for an agent to choose from - this can either be a strictly ‘legal’ set of actions, or a relaxed set of ‘permissible’ actions if no fully satisfactory options are available in the strict set. The ‘ethical’ dimension in that work is implemented as a centralised supervisor mechanism overseeing an agent during both learning and deployment, and - as discussed - the normative constraints are hard-defined, which makes this the most top-down of our case studies.
@@ 3.3 Case Study: Constitutional AI
As a second case study, we present the example of Constitutional AI, developed and implemented by ? (?). This approach is specific to the type of systems known as Large Language Models (LLMs), which were originally developed for the purpose of next-token prediction or text completion. Ethically-informed RL is used here at the fine-tuning stage, i.e., after a model has been pre-trained on a large amount of linguistic data. As such, the end product of this process is not an agent but a predictive system; however, as we discuss below, the ethical fine-tuning step of the training can be perceived as agent-based, and makes the behaviour of the LLM more goal-directed, so still has relevance to the present discussion.
@@
Constitutional AI involves a set of LLMs providing feedback to some LLM being currently trained. The constitutional methodology has been proposed as a potentially safer, more controlled and interpretable alternative to the fully bottom-up RLHF approach used in other LLMs (?). The LLMs providing feedback represent a ‘constitution’ of different principles, where each principle is defined via explicit prompts (e.g., ‘Please choose the assistant response that’s more ethical and moral. Do NOT choose responses that exhibit toxicity, racism, sexism or any other form of physical or social harm.’). The principles themselves are based on a combination of human defined preferences such as the UN Declaration of Human Rights, certain digital companies’ terms of service (to reflect the more recent digital dimensions of safety), and a set of other preferences defined by a team of researchers behind Constitutional AI (?) and others (?). The feedback from these principles is used to train a reward model for rating the outputs of the to-be-tuned LLM as ‘good’ or ‘bad’ according to its core principle.
@@
Using the feedback from this reward model as a training signal, the weights of the LLM are fine-tuned with RL. Thus, the model is fine-tuned to be more likely to produce outputs which would be considered safe by a constitution of potential ‘critic’ models with diverse preferences. An extension of this approach based on crowd-sourced constitutional principles is called Collective Constitutional AI (?) and may prove promising in the future in generating more generally or pluralistically aligned agents.
@@
In summary, according to this methodology the moral principles are explicitly defined via prompts, but the mapping between those principles and a change in behaviour (i.e., the change in language model outputs) is based on a learned representation which is very difficult for humans to interpret, thus posing a disadvantage in terms of safety. This makes Constitutional AI the most bottom-up approach of our case studies. Furthermore, while the approach is certainly promising, it must be noted that ? (?) only managed to use the constitutional principle for improving a model’s harmlessness - it remains to be seen how successful the methodology would be at promoting other values such as helpfulness.
3.4 Case Study: Reinforcement Learning in Social Dilemmas
@@
Next, we consider a case study that combines top-down principles and learning in a more equal weighting. The approach described here relies on the domain of two-player general-sum games known as social dilemmas (?, ?). We introduce these games subsequently, before describing learning agent implementations. Though these games constitute stark simplifications of reality, they are intended to model and reflect the general underlying structure of a multitude of multi-agent everyday situations.
@@ 3.4.1 Social Dilemma Matrix Games
Social dilemma games simulate situations in which each agent makes decisions facing a trade-off between individual interest and societal benefit. A widely studied subset of these games is the two-player, two-action matrix game, in which players simultaneously choose an action (e.g., Cooperate or Defect) without the opportunity to communicate. Each player then receives a certain payoff which depends on the pair of actions chosen. Higher payoff here represents the fact that an outcome is preferred by the player. What makes it a dilemma is the fact that players maximising their individual payoffs are likely to learn policies that end up being sub-optimal in terms of long-term outcomes, both for the individual and for the society of agents.
Three classic matrix dilemma games from Economics and Philosophy that are relevant to moral choice are the Prisoner’s Dilemma (?), Volunteer’s Dilemma (or ‘Chicken’; ?), and Stag Hunt (?) - payoffs for the row vs. column player are presented in Table 2. In the Prisoner’s Dilemma, mutual cooperation would achieve a Pareto-optimal outcome of $3$ points each (a desirable situation in which one player cannot be made better off without disadvantaging the other) - but each individual player’s best response is to defect due to greed (facing a cooperator, they benefit from defecting and getting the highest payoff equal to $4$) and fear (facing a defector, they suffer by cooperating and getting the lowest payoff equal to $1$). In the one-shot game, this reasoning applied by both players leads to mutual defection and thus a sub-optimal outcome, in which both players only get a payoff equal to $2$. In the Volunteer’s Dilemma, a selfish or rational player may choose to defect due to greed (i.e., not volunteer in hope that someone else does), but if both do so, both obtain the lowest possible payoffs (i.e., no one volunteers, and the society suffers). Finally, in the Stag Hunt game, two players can cooperate in hunting a stag and thus obtain the greatest possible payoff each; however, given a lack of trust between the players (i.e., each player fears a non-cooperative partner), either may be tempted to defect and hunt a hare on their own instead, decreasing both players’ payoffs as a result.
@@
Thus, in all three games the best-response (i.e., ‘rational’) strategies are not the ones leading to the better mutual outcome. ? (?) argued that repeated versions of these dilemma games in particular are relevant to morality, since, with repeated interaction, complex sets of strategies and social dynamics can evolve - for example, involving actions to punish one’s opponents for past wrongdoing, or to influence their future actions. Due to instabilities in the multi-agent environment, calculating predicted equilibria in these situations is not always computationally feasible, so simulation methods are required in order to study potential emergent behaviours and outcomes (?, ?).
Other types of games in which it is possible to model human morality are sequential games (?, ?, ?), games with more than two players (?, ?), and games with incomplete information, first introduced by ? (?). We provide these in Appendix A for context, but focus on the simpler and more abstract two-player matrix games in the subsequent formalisation due to their generalisability to a wide range of situations.
@@ 3.4.2 Reinforcement Learning in Social Dilemmas
Repeated versions of the games can be implemented as learning environments for RL agents. A social dilemma environment can be implemented for RL agents as a two-player iterated game, played over $T$ iterations (?). This is summarised in Figure 4. At each iteration, a moral player $M$ and an opponent $O$ play the one-shot matrix game corresponding to a classic social dilemma (see Table 2 for examples). At every time step $t$, the learning agent $M$ observes a state, which is the pair of actions played by $O$ and $M$ at the previous time step $t-1$: $s_{M}^{t}=(a_{O}^{t-1},a_{M}^{t-1})$, and chooses an action (Cooperate or Defect) simultaneously with the opponent $O$: $a^{t}_{M},a^{t}_{O}\in\{C,D\}$. The player $M$ then receives a reward $R_{M}^{t+1}$ and observes a new state $s_{M}^{t+1}$. In traditional RL, this reward would be provided by the environment (i.e., extrinsic reward), and in the social dilemma games it would correspond directly to the payoffs from the game (presented in Table 2).
@@ 3.4.3 Moral Preferences as Intrinsic Rewards
? (?) and ? (?) draw interesting parallels between intrinsic motivation and rewards. The implication for RL is that effective learning can result from modifying the agent such that they take into consideration an internal reward signal rather than only rewards from the environment. Early successes with this approach were reported by ? (?) in curiosity-driven exploration, where curiosity was modelled as an intrinsic reward obtained when the agent visits unfamiliar states.
@@
The intrinsic rewards framework is directly applicable to the idea of learning morality from experience. A small set of existing works in Multi-Agent RL modified the reward signal for agents in social dilemma games to embed an internal moral value or preference into reward functions (?, ?, ?, ?, ?). In these works, moral preferences have either been expressed as an operation on the extrinsic rewards (applying some function over the player’s own reward and that of their opponents), or a conditional function based on actions performed previously. A related line of work modifies reward functions to include an awareness of the opponents’ learning (?, ?) - but these are not strictly morality-focused implementations and thus lie beyond the scope of the studies reviewed here.
@@
A visualisation of the intrinsic rewards approach is presented in Figure 4. In this formulation, for player $M$, we contrast traditional agents that learn according to an extrinsic game reward $R_{M_{extr}}$, which is simply the payoff associated with the joint actions $a_{M}^{t},a_{O}^{t}$ (as defined in Table 2), versus agents that learn according to an intrinsic reward $R_{M_{intr}}$, based on some moral principle. In Section 3.4.4, we expand on the types of principles that can be encoded as intrinsic rewards.
@@
Figure 4: A step in a two-player game modelled as an RL process, from the point of view of a learning moral agent $M$ playing against a learning opponent $O$. Agent $M$ calculates their moral (intrinsic) reward according to their own ethical framework, which may or may not consider the game (extrinsic) reward coming from playing the dilemma.
@@
In practice, both extrinsic and intrinsic rewards can be combined in a multi-objective manner (?) to create an agent able to pursue multiple goals (for example, trying to obtain both moral reward and the reward from the task at hand - ?). This was the approach used by ? (?) to create their ‘curious’ yet task-specific agent. Furthermore, multi-objective RL approaches can also be used to combine multiple intrinsic moral rewards to create a multi-faceted moral agent (e.g., see the multi-objective moral agent from ?), which allows for the creation of more complex and comprehensive moral agents.
@@
This case study implements a decentralised RL approach where each agent chooses an action in the game without a central planner. A parallel line of work exists in the modelling and ethics literature, where an ‘ethical’ agent is implemented as a centralised mechanism on the environment - this has applications in entirely different domains such as governmental policy design (?), and lies beyond the scope of this work, which focuses on the moral alignment of individual agents.
3.4.4 Formalising Existing Ethical Frameworks as Rewards for AI Agents
@@
A variety of moral frameworks has been developed over the millennia in the fields of Moral Philosophy, Psychology, Biology and Economics. We review a representative set of these in Table 3, formulating them as potential reward functions for agents in two-player games. A growing body of RL work demonstrates that mathematical formulations of these frameworks are able to effectively encourage agents to learn more prosocial and cooperative behaviour in social settings (?, ?, ?, ?).
@@
Utilitarian or Altruistic ?, ?, ?, ?, ? Sum of all players’ payoffs, with various weights on on own vs. others’ payoffs. either Consequentialist
@@
Deontological ?, ?, ?, ?, ? Positive (or negative) reward if the action taken adheres to (or violates) a pre-defined norm. Agent may enforce norm in others or not. External Norm-based
@@
Virtue Ethics (can be based on other types above) ?, ?, ? Reward for adhering to a single specific virtue (e.g., cooperativeness), or a mix of virtues (e.g., sum cooperativeness + equality rewards). Internal either or both
@@
Table 3: Moral frameworks proposed in past literature in Philosophy, Psychology, Biology and Economics, and potential implementations in terms of rewards for RL agents.
@@
We classify possible moral agent types in Table 3 into Consequentialist or Norm-based. Consequentialist morality focuses on the outcomes or consequences of an action. In dilemma games, this can be easily implemented as an operation on the two players’ payoffs, including objective functions that are Utilitarian (?), defining actions as moral if they maximise total utility for all agents in a society, or Altruistic (?), where an agent is inclined to maximise their opponent’s winnings, perhaps without even considering their own. Consequentialist morality also includes fairness-based objective functions. These can be based on the difference between two players’ rewards, including the inequity aversion models of ? (?, ?), or other calculations of the ratio between payoffs with different preferences for fair or equal outcomes (?, ?, ?). Finally, consequentialism also accounts for the principles of distributive justice of ? (?), in which agents distribute resources without knowing where they will end up in that distribution.
@@
Norm-based morality, on the other hand, focuses on defining actions as moral so long as they adhere to a society’s external norms or to a specific agent’s duties, regardless of the consequences further down the line. This framework includes the Deontological ethics of ? (?), which identifies specific acts as morally required, permitted or forbidden. An example norm for dilemma game-playing agents may be conditional cooperation, in which an agent is expected to cooperate against a cooperative opponent (?). Further norm-based objective functions have been formalised by ? (?, ?, ?) and ? (?), who propose various weights on personal versus social importance of the norm at hand.
@@
A third type of ethical framework that is often distinguished is Virtue Ethics of ? (?) - in this line of reasoning, moral agents act according to their certain internal virtues. In practice, these virtues themselves often have Consequentialist or Norm-based foundations - consider, for example, the virtues of fairness or care, respectively (?, ?). The key distinction of Virtue Ethics is that a single agent may rely on more than one type of virtue, and different agents may weigh the virtues differently against one another (?, ?), so a more expressive way of modelling virtue ethics might be through a weighted, multi-objective paradigm - such as that suggested by ? (?).
@@
In Table 3, we additionally distinguish agents by the external or internal nature of their moral motivation - while each of these frameworks can be formalised as an intrinsic reward function for the RL agent, sometimes the moral aspect of the intrinsic motivation comes from outside the agent themselves. An example of this is where certain moral norms exist externally in a society, so the agent prefers to adhere to them in their individual behaviour.
@@
The final step for formalising morality as intrinsic motivation for RL agents, then, is to formulate precise mathematical definitions of rewards based on these frameworks. Using the domain of social dilemmas introduced above, ? (?) propose a formulation in terms of actions and/or outcomes at some time step $t$ in a two-player, two-action social dilemma game. For illustration, we present their mathematical definitions of five types of moral rewards in Table 4 (?).
@@
Table 4: Definitions of a sub-set of intrinsic moral reward types, from the point of view of the moral agent $M$ playing a social dilemma game versus an opponent $O$.
@@
Norm-based morality, on the other hand, is operationalised in terms of a positive reward for an agent whose action $a^{t}_{M}$ adheres to a given norm, or a negative reward for actions violating a given norm. For example, a Deontological agent following the norm of conditional cooperation (defined above) receives a negative reward of value $\xi$ for defecting against an opponent who previously cooperated. The other norm-based agent Virtue-kindness follows a simple cooperative norm, obtaining a positive reward $\xi$ for cooperating, regardless of what their opponent did previously.
Finally, these examples also outline one multi-objective reward function Virtue-mixed, which linearly combines the two separate virtues ‘equality’ and ‘kindness’, with a weighting parameter $\beta$, and a parameter $\hat{\xi}$ used to fit the ‘kindness’ element of the reward to the same scale as equality.
? (?) implement the five moral intrinsic rewards defined in Table 4 (as well as a Selfish baseline agent) and show that Q-learning agents learning against a variety of opponents are able to learn optimal moral policies given sufficient exploration. A further study by ? (?) extends this analysis to populations involving a partner selection mechanism and a larger amount of possible prosocial and antisocial moral agent types.
3.5 Case Study: Fine-tuning LLM Agents with Intrinsic Rewards
Finally, to complement the intrinsically moral RL agents described above, we consider the case of fine-tuning LLM-based agents for moral alignment. This work relies on the use of LLMs as decision-makers (?, ?, ?) in strategic scenarios (?, ?, ?), often studied using game-theoretic environments such as social dilemma games discussed above (?, ?). Thus, the approach of fine-tuning LLMs with RL reconciles the previous implementations of hybrid moral alignment in pure-RL agents with the more general and currently popular domain of LLM agents.
As discussed previously, the predominant approach for aligning LLMs to human values is by training reward models on human preference data (i.e., RLHF), and fine-tuning an LLM using that reward model, though non-RL approaches, such as DPO (?), enable preference-based fine-tuning without the need for RL-based training. This approach teaches the LLM the values that are implicit in the human preference data, without any explicit definition of what they are. A core disadvantage of this, as discussed in Section 2, is the lack of transparency interpretability and control, and the risks of reward-hacking.
? (?) instead propose a methodology for fine-tuning LLMs based on intrinsic rewards rather than human-generated preference data (as in the case of RLHF). The intrinsic rewards in this approach are defined in terms of actions or consequences in a specific environment, but the policies learned via such fine-tuning can, in theory, generalise to other environments. ? (?) present a study fine-tuning LLM agents using the rich domain of social dilemmas described in Section 3.4. For LLM agents in particular, decision-making scenarios or games can be expressed in the form of text, with specific tokens corresponding to each action choice. During fine-tuning, the model learns to associate certain action tokens and strategies with rewards, with online RL implemented via on-policy methods such as PPO (?, though offline implementations are also possible, see ?).
The results in ? (?) offer some evidence of the ability of LLM agents to generalise moral policies learned in one game to other environments characterised by a similar game structure - specifically, from the IPD agent to four other 2x2 matrix games. The challenges in this implementation include the fact that fine-tuning on a small set of specific action tokens may make the models over-fit to those tokens - indeed, ? (?) observed that the fine-tuned model was more likely to produce the action tokens in response to unrelated questions of a similar format. Nevertheless, overall, this work suggests an interesting emerging direction in using intrinsic rewards as a hybrid and more controllable and interpretable implementation of moral alignment in agentic LLM systems.